MILLIMETER WAVE SPECTROSCOPY OF THE INTERMOLECULAR STRETCHING BAND OF He-HCN.

\underline{KENSUKE HARADA}, KEIICHI TANAKA, and TAKEHIKO TANAKA, Department of Chemistry, Faculty of Sciences, Kyushu University, Hakozaki, Higashiku, Fukuoka, 812-8581 JAPAN; SHINKO NANBU, Institute for Molecular Science, Okazaki, 444-8585, Japan; MUTSUMI AOYAGI, Computing and Communications Center, Kyushu University, Hakozaki, Higashiku, Fukuoka, 812-8581 JAPAN.

Millimeter-wave absorption spectroscopy combined with a pulsed-jet expansion technique was applied to the observation of the intermolecular stretching excited state (ν_{s}) and internal rotation second excited state (j = 2) of the He-HCN complex. Two supersonic jet nozzles and 10 round trip multi-reflection optical path were used for the measurement. Transitions belonging to the ν_{s}, ν_{s} ← j = 1, j = 2 ← 1, and j = 2 ← 0 bands were observed in the frequency region of 150–280 GHz, where ν_{s} refers to the intermolecular stretching excited state, and j = 1 and 2 to internal rotation excited states. These transitions were definitely assigned using combination differences. The observed transition frequencies as well as the transition frequencies of the internal rotation fundamental band^{a} (j = 1 ← 0) were analyzed to improve an empirical intermolecular potential energy surface. The intermolecular stretching frequency obtained is 9.1618 cm^{−1}, while the dissociation energy (D_{0}) is 9.4411 cm^{−1}. The average distance < R > from the He atom to the center of mass of HCN and the root mean square amplitude of the intermolecular stretching vibration is 5.572 Å and 2.182 Å for the ν_{s}, J = 1 state, which is far longer and larger than those for the ground state (4.306 Å and 0.576 Å ). The wavefunction of the intermolecular stretching excited state has a node along the radial coordinate and the probability density extends up to more than 12 Å .

aK. Harada, K. Tanaka, T. Tanaka, S. Nanbu, and M. Aoyagi J. Chem. Phys. \underline{117}, 7041, (2002).